# How connection probability shapes fluctuations of neural population dynamics

### v1

@@ IV.3 Linear response of the mean activity
Figure 5: Linear response of the time-dependent mean population rate (“rate susceptibility”) to a weak modulation of the external current. (a) $\tilde{\chi}_{r}$ for the microscopic model with quenched (green circle) and annealed (red squares) networks for $p=0.1$, $d=0$ ms, and the corresponding 1st- (orange) and 2nd-order (blue) mean-field theories, Eq. (35) (where $p$ is set to $1$ for the 1st-order MF theory). Top and bottom row: amplitude $|\tilde{\chi}_{r}|$ and phase $\phi_{r}$ of the susceptibility, respectively. Dashed horizontal line: $\phi_{r}=-\pi/2$. (b) Same for $p=0.4$, $d=0.001$ s. Other parameters in (a) and (b): $N=1000$, $w=-1$ mVs.
We now analyse how the network responds to a small, time-dependent perturbation of the external current, $\mu(t)=\mu_{0}+\mu_{1}(t)$ by calculating the linear response function [67]. To this end, we linearize our mesoscopic mean-field model, Eq. (20), around the fixed point and obtain a generalization of Eq. (30):
@@
[…]
@@
The rate susceptibility is in general complex-valued with absolute value $|\tilde{\chi}_{r}(\omega)|$ and phase $\phi_{r}(\omega)$. A straightforward interpretation is that the system driven by a sinusoidal external input modulation $\mu_{1}(t)=\epsilon\sin(\omega t)$ with small amplitude $\epsilon$ will respond with a sinusoidal rate modulation with amplitude $\epsilon|\tilde{\chi}_{r}(\omega)|$ and a phase shift $\phi_{r}(\omega)$. For all models studied here, the amplitude decreases for high frequencies with a power law with exponent -1 (Fig. 5). The annealed network and the 2nd-order MF model agree very well, in contrast to the 1st-order MF model which massively overestimates the amplitude at high frequencies. The quenched network shows a slightly smaller amplitude than predicted by the 2nd-order MF model, but is still reasonably well approximated by the latter. While the phase shift vanishes at low frequencies as expected, the phase shift approaches $-\frac{\pi}{2}$ in the limit of large frequencies (Fig. 5). We can understand this behaviour analytically by considering the limits of Eq. (35). For $\omega=0$ we obtain
@@
which is real-valued and has therefore a phase zero. The amplitude in the low-frequency limit is approximately the negative inverse of the coupling strength $-w^{-1}$, under the condition that $1\ll wF_{h}$ and $C$ large. Similarly, for large frequencies we find asymptotically
@@
In this limit, the phase converges to $-\frac{\pi}{2}$ and the amplitude follows a power law $\omega^{-1}$ scaled by $F_{h}/\tau$. The different values for the derivative $F_{h}$ explain the difference between the 1st-order and the 2nd-order MF model. To understand this difference analytically, we use the following approximation for the slope of the transfer function at the fixed point (see Appendix D):
@@
defined on the interval $[0,1]$. This function has an inverted-U shape and is symmetric with respect to $x=\frac{1}{2}$ (a graph of $Q$ is shown in Fig. 8c, dotted lines). Equation (38) clearly shows that the factor $F_{h}$ increases with the connection probability $p$ because $\sigma_{0}^{2}$ decreases with $p$, cf. Eq. (27). The first-order MF theory corresponding to the mean-connectivity network (without additional noise) is obtained by setting $p=1$ in Eq. (27). In this case, the dispersion $\sigma_{0}^{2}$ vanishes, and hence the slope $F_{h}$ is maximized. Therefore, the 1st- and 2nd-order models predict the same power-law behaviour, but with a higher magnitude for the 1st-order model (Fig. 5a). With increasing connection probability $p$ the susceptibility approaches the 1st-order MF theory. The low-frequency limit is in both mesoscopic models almost independent of $p$ as the low-frequency limit is roughly given by $-w^{-1}$. Simulations of the annealed network model match the predictions of the 2nd-order MF model while the quenched network model has slightly lower amplitude and slightly lower phase for intermediate frequencies. Overall, the 2nd-order MF theory yields a strongly improved prediction compared to the 1st-order MF model.
The synaptic delay $d$ appears in $\tilde{\chi}_{r}$ only via the complex exponential in Eq. (34). On the other hand, the delay does not influence the low-frequency limit and the high-frequency asymptotics, Eqs. (36) and (37) of the response, respectively. For intermediate frequencies the complex exponential causes a modulation and a peak in the amplitude as a function of $\omega$ (Fig. 5b). Simulations of the annealed network model perfectly confirm the prediction of the 2nd-order MF model, whereas the quenched network model has a less pronounced peak at a lower frequency.
@@ IV.4 Power spectral density
Figure 6: The first-order mean-field theory strongly underestimates the power spectral density of the population activity. Simulations of the power spectral densities of the population activity $A_{N}(t;\Delta t)$ for the microscopic (green circles: quenched network, red squares: annealed network) and mesoscopic models (orange triangle: 1st-order MF model, blue diamonds: 2nd-order MF model). Solid lines depict the 1st- and 2nd-order MF theory, Eqs. (59) and (42), respectively. Parameters: $w=-1$ mVs a) $N=500$, $p=0.07$, $\mu(t)=\mu_{0}=10$ mV.
Now we turn to the second-order statistics of the stationary dynamics. We assume that the external stimulus $\mu(t)$ is a stationary process with given mean $\mu_{0}$ and power spectral density $S_{\mu\mu}(\omega)$. The second-order statistics can be characterized, in the frequency domain, by the power spectral density matrix of the mean-field state vector $\boldsymbol{X}(t)$ (cf. Eq. (30)). This matrix involves, e.g., the power spectrum of the average input potential $\bar{h}(t)=\bar{h}_{0}+X_{1}(t)$ but also cross spectra such as between $\bar{h}(t)$ and $\sigma^{2}(t)=\sigma_{0}^{2}+X_{2}(t)$. Furthermore, we are interested in the power spectra of the average intensity $r(t)$ and the population activity $A(t)$, as well as in the variance of $r(t)$. The latter is studied in the subsequent section.
The power spectral density matrix $\operatorname{S}(\omega)$ is defined via the relation
@@
The matrix $\operatorname{S}$ gives us the (cross-) spectral densities between $\bar{h}$, $\sigma^{2}$ and $\xi$. For the power spectral densities of the stochastic firing rate $r(t)$ and the activity $A(t)$, we use the linear approximation
$\tilde{r}\approx F_{h}\tilde{h}+F_{\sigma}\widetilde{\sigma^{2}}+\frac{1}{\sqrt{N}}\tilde{\xi},\quad\tilde{A}\approx\tilde{r}+\sqrt{\frac{r_{0}}{N}}\tilde{\eta}.$ (41)
These linear equations allow us to compute the power spectral density $S_{rr}(\omega)$ of the stochastic population rate $r(t)$ via the relation $\left\langle\tilde{r}(\omega)\tilde{r}^{*}(\omega^{\prime})\right\rangle=2\pi S_{rr}(\omega)\delta(\omega-\omega^{\prime})$, which yields
$S_{rr}(\omega)\approx F_{h}^{2}S_{11}(\omega)+F_{\sigma}^{2}S_{22}(\omega)+\frac{1}{N}S_{33}(\omega)+\\ 2F_{h}F_{\sigma}\mathfrak{Re}(S_{12}(\omega))+2\frac{F_{h}\mathfrak{Re}(S_{13}(\omega))+F_{\sigma}\mathfrak{Re}(S_{23}(\omega))}{\sqrt{N}}.$
Further, we find, using Eq. (41), the power spectrum of $A(t)$ within the linear approximation as
@@
In Fig. 6, we compare the theory for the power spectrum of $A(t)$, Eq. (42), to simulations of $A_{N}(t,\Delta t)$ in the microscopic and mesoscopic models. For all models, the power spectra exhibit a trough at low frequencies. The 1st-order MF model massively underestimates the observed low-frequency power, whereas the 2nd-order MF model matches well the microscopic model with annealed connectivity. Compared to the quenched network, the low-frequency power of the annealed network is still significantly reduced. Thus, despite the considerable improvement compared to the 1st-order MF theory, the underestimation of low-frequency power reveals a short-coming of the annealed approximation.
@@ IV.5 Variances
Figure 7: Stationary variance of the population rate and mean input potential for varying connection probability $p$. (a) The number of neurons is fixed to $N=1000$ (hence $C=1000p$). (b) The in-degree is fixed to $C=100$ (hence $N=100/p$). Left panels of (a) and (b): Variance of the population rate $r(t)$ (Symbols: simulations of microscopic networks and the mesoscopic MF models as indicated in the legend; orange and blue solid lines: full theory Eq. (43) and Eqs. (62); black dashed line: analytical approximation for the variances based on the 2nd-order MF theory, Eq. (45), (46) and Eqs. (44), (48) for $\text{var}(r)$ and $\text{var}(\bar{h})$, respectively). Right panels of (a) and (b): Same as left panels but for variance of $\bar{h}$. Parameters: $\mu_{0}=10$ mV, $w=-1$ mVs, $\sigma_{\text{ext}}=0$ mV, $\beta=5$ mV-1
@@
where $\sigma_{ij}=\left\langle X_{i}(t)X_{j}(t)\right\rangle$ are the elements of the covariance matrix of the state vector $\boldsymbol{X}(t)$. In particular, $\sigma_{11}$ yields the variance of the population-averaged input potential $\bar{h}(t)$. Approximate analytical expressions for $\sigma_{ij}$ can be calculated using the linearized system, Eq. (32), and are given in the appendix, Sec. G. Our theory, Eq. (43), agrees well with microscopic simulations (Fig. 7 and 8), discussed in more detail below. In general, we find a good quantitative match between the theory and simulations of the annealed network. While, in the case of quenched random connectivity, the theory again exhibits quantitative deviations, our 2nd-order MF theory still captures the parameter dependence of the variance qualitatively and yields a much better quantitative prediction than the 1st-order MF theory.
@@ IV.5.1 Dependence on the connection probability
Which connection probability causes the largest variability in our network model? To address this question, we analyze the variances of the population rate $r$ and the average input potential $\bar{h}$ as a function of $p$ (Fig. 7). To avoid a trivial dependence through the mean firing rate, we vary $p$ such that the mean firing rates remain roughly unchanged. According to Eq. (26), a constant rate can be achieved by keeping the coupling strength $w=NpJ$ constant. In an experimental setting, where the synaptic efficacy $J$ is the basic physiological parameter, this constraint could be realized in two different scenarios, either by fixing the network size $N$ and changing the synaptic efficacy $J$ such that $Jp=\text{const.}$., or by fixing both the number of input connections $C=Np$ and the synaptic efficacy $J$.
@@
In the first scenario with constant $N$, we find that the variance of the rate $r(t)$ increases with connection probability in microscopic network simulations (Fig. 7a, left), although the variance of the input potential $\bar{h}$ decreases (Fig. 7a, right). The 1st-order MF model cannot explain this behavior because the model equation (21) does not depend on the parameter $p$ when $w$ is fixed and $N$ is constant. Thus, for all $p\in[0,1]$, the variance predicted by the 1st-order MF model corresponds to the fully connected case $p=1$, i.e. to the mean-connectivity network. In contrast, the dependence on $p$ of the variances measured from simulations of the annealed network are quantitatively well reproduced by the 2nd-order MF theory, Eq. (43). The 2nd-order theory also captures the variances of the quenched network qualitatively. The quantitative agreement is also reasonably good, however, we find that the rate variance is overestimated and the variance of the population-averaged input potentials is underestimated by the annealed network. Nevertheless, the prediction of variances by the 2nd-order theory yields a significant improvement over the 1st-order MF theory.
In order to gain an analytical understanding, we use a simplified theory based on an approximation of the variance of $\bar{h}$ when $-wF_{h}\gg 1$ (see Appendix, Sec. G) :
$\text{var}(\bar{h})\equiv\sigma_{11}\approx\left(\frac{\mu_{0}}{2\tau N}+\frac{\sigma_{\text{ext}}^{2}}{-2w}\right)\frac{1}{F_{h}}.$ (44)
Furthermore, for the parameters used in this study, we find that the variance of $r$ is vastly dominated by the first term in Eq. (43), while the other terms are small enough to be ignored. With Eq.(44), we thus obtain the following approximation
@@
The variance formulas (44) and (45) share a common factor that represents the sum of two contributions to the variability. The first term of the sum corresponds to the intrinsically generated variability caused by finite-size fluctuations of order $\sqrt{r_{0}/N}$ in the mesoscopic MF model, Eq. (20). We recall that the dependence of these fluctuations on the firing rate $r_{0}=-\mu_{0}/w$, arises from the Poisson spiking noise at the microscopic level. Thus, the Poisson property that the variance scales proportionally to the mean causes finite-size fluctuations that increase with the mean stimulus strength $\mu_{0}$ (see below). We emphasize that this contribution to the variability is a clear finite-size effect that vanishes in the limit $N\to\infty$. In contrast, the second term of the sum in Eqs. (44) and (45) represents externally generated variability proportional to the variance of the common external stimulus.
Figure 8: Variance of the stochastic population rate $r(t)$ for varying mean external stimulus $\mu_{0}$. (a) Moderate-size network of $N=10^{3}$ neurons whose inputs exhibit internally generated, finite-size noise but no external noise, $\sigma_{\text{ext}}=0$. Left: Response to a step stimulus, where $\mu_{0}$ jumps from $28$ mV to $50$ mV at time $t=0.04$ s. Blue and orange line: simulations of the 2nd-order MF model for $C=100$ and $C=1000$, respectively (the latter case ($p=1$) corresponds to the 1st-order MF model or mean-connectivity approximation of the former case ($p=0.1$)); black dotted lines: theoretical prediction of stationary variances, Eq. (43); gray dashed line: stimulus onset. The inset magnifies the same data for $C=100$. Right: Stationary variance of $r(t)$ vs. mean stimulus strength $\mu_{0}$. (Symbols: simulations of quenched network (green circles occluded by red squares), annealed network (red squares), both for $C=100$, 1st- (yellow triangles) and 2nd-order MF model (blue diamonds). Solid lines: full 1st- (yellow) and 2nd-order MF theory (blue); black dashed lines: corresponding analytical approximations, Eq. (46).) The gray dashed line marks the border of the region of interest $\mu_{0}\leq-wr_{\text{m}}/2$, where the hazard function at the fixed point, $\phi(\bar{h}_{0})$, is convex. (b) Analogous to (a) but for a much larger network ($N=5\cdot 10^{4}$) with negligible, internally generated, finite-size noise but additional external noise, $\sigma_{\text{ext}}^{2}=1$ mV2. The connection probability is varied by different values of $C$ as indicated. The yellow line corresponds to the common 1st-order MF model or mean-connectivity network. (c) Slope $\hat{F}_{h}$ vs. mean stimulus $\hat{\mu}_{0}$. Dotted line: $p=1$ corresponding to the function $Q$, solid line: $p=0.7$. Gray dashed line as in (a),(b) right. (d) Same for the product $\hat{\mu}_{0}\hat{F}_{h}$.
@@
In the second scenario with constant $C$ and varying population size $N=C/p$, we observe different dependencies of the variances on $p$ compared to the first scenario with constant $N$ (Fig. 7b). In simulations of microscopic networks and 2nd-order MF model, the rate variance increases supralinearly with $p$ and the variance of the mean input potential $\bar{h}$ exhibits a non-monotonic behavior with a maximum inside the range $10<p<$ 1. Again, the 1st-order MF theory does not capture these behaviors. Note that there is a good qualitative agreement between 2nd-order MF theory and microscopic network simulations. As in the first scenario, however, the variance of the average input potential in the quenched network is underestimated by the annealed network, which, in turn, is quantitatively well matched by the 2nd-order MF theory.
@@
(cf. Eq. (38)). The expression for the rate variance in Eq. (46) is a product of two monotonously increasing functions of $p$ – a linear function (the prefactor of $\hat{F}_{h}$) and a strictly convex function ($\hat{F}_{h}$). Hence, the variance of the population rate is a monotonosly increasing, strictly convex function of $p$ explaining our observation in Fig. 7b (left). In contrast, the variance of the mean input potential, Eq. (44), can be rewritten for constant $C$ and dimensionless parameters as

### v2

@@ IV.3 Linear response of the mean activity
We now analyze how the network responds to a small, time-dependent perturbation of the external current, $\mu(t)=\mu_{0}+\mu_{1}(t)$ by calculating the linear response function [74]. To this end, we linearize our mesoscopic mean-field dynamics, Eq. (20), around the fixed point and obtain a generalization of Eq. (29):
@@
Here, $\operatorname{I}_{3}$ denotes the $3\times 3$ identity matrix and the tilde notation $\tilde{f}(\omega)$ for a function $f(t)$ denotes its Fourier transform, i.e. $\tilde{f}(\omega)=\int_{-\infty}^{\infty}f(t)e^{-i\omega t}\,dt$. The explicit expressions for the elements of $\tilde{\operatorname{\chi}}$ are given in the appendix, Eq. (63).
@@
The rate susceptibility is in general complex-valued with absolute value $|\tilde{\chi}_{r}(\omega)|$ and phase $\phi_{r}(\omega)$. A straightforward interpretation is that the system driven by a sinusoidal external input modulation $\mu_{1}(t)=\epsilon\sin(\omega t)$ with small amplitude $\epsilon$ will respond with a sinusoidal rate modulation with amplitude $\epsilon|\tilde{\chi}_{r}(\omega)|$ and a phase shift $\phi_{r}(\omega)$. For all models studied here, the amplitude decreases for high frequencies with a power law with exponent -1 (Fig. 3d-f). The MF2 theory agrees very well with the quenched microscopic network, in contrast to the MF1 theory which massively overestimates the amplitude at high frequencies. The quenched network shows a slightly smaller amplitude than predicted by the 2nd-order MF theory, but is still reasonably well approximated by the latter. While the amplitude vanishes at low frequencies as expected, the phase shift approaches $\frac{\pi}{2}$ in the limit of large frequencies (Fig. 3d-f). We can understand this behavior analytically by considering the limits of Eq. (34). For $\omega=0$ we obtain
@@
which is real-valued and has therefore a phase zero. The amplitude in the low-frequency limit is approximately the negative inverse of the coupling strength $w^{-1}$, under the condition that $1\ll wF_{h}$ and $C$ large. Similarly, for large frequencies we find asymptotically
@@
In this limit, the phase converges to $\frac{\pi}{2}$ and the amplitude follows a power law $\omega^{-1}$ scaled by $F_{h}/\tau$. The different values for the derivative $F_{h}$ explain the difference between the MF1 and the MF2 theory. To understand this difference analytically, we use the following approximation for the slope of the transfer function at the fixed point (see Appendix D):
@@
defined on the interval $[0,1]$. This function has an inverted-U shape and is symmetric with respect to $x=\frac{1}{2}$ (a graph of $Q$ is shown in Fig. 6c, dotted lines). Equation (37) clearly shows that the factor $F_{h}$ increases with the connection probability $p$ because $\sigma_{0}^{2}$ decreases with $p$, cf. Eq. (28). The MF1 theory corresponding to the mean-connectivity network (without additional noise) is obtained by setting $p=1$ in Eq. (28). In this case, the dispersion $\sigma_{0}^{2}$ vanishes, and hence the slope $F_{h}$ is maximized. Therefore, the MF1 and MF2 theories predict the same power-law behavior of the rate susceptibility with $\omega$ at large frequencies, but with a higher magnitude for the 1st-order model (Fig. 3d). With increasing connection probability $p$ the susceptibility approaches the 1st-order MF theory. The low-frequency limit is in both mesoscopic models almost independent of $p$ as the low-frequency limit is roughly given by $w^{-1}$. Simulations of the annealed network network (not shown in the figure) match the predictions of the MF2 theory while the quenched network model has slightly lower amplitude and slightly lower phase for intermediate frequencies. Overall, the MF2 theory yields a strongly improved prediction compared to the MF1 theory.
The synaptic delay $d$ appears in $\tilde{\chi}_{r}$ only via the complex exponential in Eq. (33). On the other hand, the delay does not influence the low-frequency limit and the high-frequency asymptotics, Eqs. (35) and (36) of the response, respectively. For intermediate frequencies the complex exponential causes a modulation and a peak in the amplitude as a function of $\omega$ (Fig. 3e). Simulations of the annealed network model perfectly confirm (not shown in figure) the prediction of the 2nd-order MF theory, whereas the quenched network model has a less pronounced peak at a lower frequency. The susceptibility of MFsparselim fits well with the quenched network for small or medium connection probability, $p=0.1$ and $p=0.4$ (Fig. 3d, e). For larger $p$, however, there is a noticeable discrepancy for MFsparselim, while MF1 and MF2 have a good match with the quenched network (at $p=0.95$, see Fig. 3f). In marked contrast to MF1 and MFsparselim, the predictions of MF2 are good over the entire range of connection probabilities $p$.
@@ IV.4 Power spectral density
Figure 4: Power spectral densities of the population activity and population rate. a) Power spectrum of $A_{N}(t;\Delta t)$ for quenched (green circles) and annealed network (red squares) and 1st- and 2nd-order theory (orange and blue lines, Eqs. (65) and (42), respectively). b) Same for the power spectrum of $r(t)$ (theory: Eq. (41)) Parameters: w=1w=-1 mVs a) $N=500$, $p=0.07$, $\mu(t)=\mu_{0}=10$ mV, $\sigma_{\text{ext}}=0$, $\Delta t=0.1$ ms.
Now we turn to the second-order statistics of the stationary dynamics. We assume that the external stimulus $\mu(t)$ is a stationary process with given mean $\mu_{0}$ and power spectral density $S_{\mu\mu}(\omega)$. We begin with the power spectrum of the population activity $A(t)$ and population rate $r(t)$. This statistics has been calculated analytically using linear-response theory for finite-size networks with uniform full connectivity [71, 47]. In a similar fashion, we here derive the power spectrum via a linearization of our 2nd-order MF dynamics. To this end, we start with the power spectral density matrix $\operatorname{S}(\omega)$ of the mean-field state vector $\boldsymbol{X}(t)$ (cf. Eq. (29)) defined via the relation
@@
The matrix $\operatorname{S}$ gives us the (cross-) spectral densities between $\bar{h}$, $\sigma^{2}$ and $\xi$. For the power spectral densities of the population activity $A(t)$, we use the linear approximation
$\tilde{A}\approx\tilde{r}+\sqrt{\frac{r_{0}}{N}}\tilde{\eta},\quad\tilde{r}\approx F_{h}\tilde{h}+F_{\sigma}\widetilde{\sigma^{2}}+\frac{1}{\sqrt{N}}\tilde{\xi}$. (40)
which yields approximate expressions for the power spectrum of the stochastic population rate $r(t)$,
$S_{rr}(\omega)\approx F_{h}^{2}S_{11}(\omega)+F_{\sigma}^{2}S_{22}(\omega)+\frac{1}{N}S_{33}(\omega)+\\ 2F_{h}F_{\sigma}\mathfrak{Re}(S_{12}(\omega))+2\frac{F_{h}\mathfrak{Re}(S_{13}(\omega))+F_{\sigma}\mathfrak{Re}(S_{23}(\omega))}{\sqrt{N}},$ (41)
and of the population activity $A(t)$:
@@
In Fig. 4, we compare the theoretical predictions, Eqs. (41) and (42), to the corresponding power spectra obtained from simulations of the quenched and annealed network. In all cases, the power spectra of the population activity display a pronounced trough at low frequencies (Fig. 4a). The MF1 theory severely underestimates the observed low-frequency power, whereas the MF2 theory matches well the microscopic model under the annealed approximation. Nonetheless, even with this marked improvement over MF1, the annealed network’s low‐frequency power remains lower than that of the quenched network, exposing a limitation of the annealed approximation. The discrepancy of the power spectra of quenched and annealed networks is caused by the partial elimination of the effect of the single-neuron auto-correlations on the recurrent input when applying the annealing (i.e. the permanent resampling of connectivity). We note that the elimination of temporal correlations only concerns the incoherent part of the recurrent input, as expressed by the white noise in Eq. (12). In contrast, for finite $N$, the coherent part of the recurrent input fluctuations (finite-size fluctuations of population activity) still contains effects of spike-train auto-correlations, even in the annealed network. Finally, we mention that we have not included MFsparselim in Fig. 4 because its stationary dynamics is characterized by a deterministic equilibrium without any fluctuations and a vanishing power spectrum.
@@ IV.5 Variances
Figure 5: Stationary variance of the population rate and mean input potential for varying connection probability $p$. (a) The number of neurons is fixed to $N=1000$ (hence $C=1000p$). (b) The in-degree is fixed to $C=100$ (hence $N=100/p$). Left panels of (a) and (b): Variance of the population rate $r(t)$ (Symbols: simulations of microscopic networks as indicated in the legend; orange and blue solid lines: full theory Eq. (43) and Eqs. (68); black dashed line: analytical approximation for the variances based on the 2nd-order MF theory, Eq. (46), (47) and Eqs. (45), (49) for $\text{var}(r)$ and $\text{var}(\bar{h})$, respectively; dash-dotted line: MFsparselim.) Right panels of (a) and (b): Same as left panels but for variance of $\bar{h}$. Parameters: $\mu_{0}=10$ mV, w=1w=-1 mVs, $\sigma_{\text{ext}}=0$ mV, $\beta=5$ mV-1
@@
where $\sigma_{ij}=\left\langle X_{i}(t)X_{j}(t)\right\rangle$ are the elements of the covariance matrix of the state vector $\boldsymbol{X}(t)$. In particular, $\sigma_{11}$ yields the variance of the population-averaged input potential $\bar{h}(t)$. Approximate analytical expressions for $\sigma_{ij}$ can be calculated using the linearized system, Eq. (31), and are given in the appendix, Sec. G. Our theory, Eq. (43), agrees well with microscopic simulations (Fig. 5 and 6), discussed in more detail below. In general, we find a good quantitative match between the theory and simulations of the annealed network. While, in the case of quenched random connectivity, the theory again exhibits quantitative deviations, our 2nd-order MF theory still captures the parameter dependence of the variance qualitatively and yields a much better quantitative prediction than the 1st-order MF theory. Again, MFsparselim without any noise in the external input deterministically converges into an equilibrium with zero variance.
@@ IV.5.1 Dependence on the connection probability
Which connection probability causes the largest variability in our network model? To address this question, we analyze the variances of the population rate $r$ and the average input potential $\bar{h}$ as a function of $p$ (Fig. 5). To avoid a trivial dependence through the mean firing rate, we vary $p$ such that the mean firing rates remain roughly unchanged. According to Eq. (26), a constant rate can be achieved by keeping the coupling strength $w=NpJ$ constant. In an experimental setting, where the synaptic efficacy $J$ is the basic physiological parameter, this constraint could be realized in two different scenarios, either by fixing the network size $N$ and changing the synaptic efficacy $J$ such that $Jp=\text{const.}$, or by fixing both the number of input connections $C=Np$ and the synaptic efficacy $J$.
@@
In the first scenario with constant $N$, we find that the variance of the rate $r(t)$ increases with connection probability in microscopic network simulations (Fig. 5a, left), although the variance of the input potential $\bar{h}$ decreases (Fig. 5a, right). MF1 cannot explain this behavior because the model equation (21) does not depend on the parameter $p$ when $w$ is fixed and $N$ is constant. Thus, for all $p\in[0,1]$, the variance predicted by the 1st-order MF theory corresponds to the fully connected case $p=1$, i.e. to the mean-connectivity network. In contrast, the dependence on $p$ of the variances measured from simulations of the annealed network are quantitatively well reproduced by the 2nd-order MF theory, Eq. (43). The 2nd-order theory also captures the variances of the quenched network qualitatively. The quantitative agreement is also reasonably good, however, we find that the rate variance is overestimated and the variance of the population-averaged input potentials is underestimated by the annealed network. Nevertheless, the prediction of variances by the 2nd-order theory yields a significant improvement over the 1st-order MF theory.
In order to gain an analytical understanding, we use the following explicit approximations of the variances (see Appendix, Sec. G):
$\text{var}(\bar{h})\equiv\sigma_{11}\approx\frac{\frac{-w\mu_{0}}{2\tau N}+\frac{\sigma_{\text{ext}}^{2}}{2}}{(1-wF_{h})},\quad\text{var}(r)\approx F_{h}^{2}\text{var}(\bar{h}).$ (44)
In the last approximation, we have used Eq. (43) and the fact that, for the parameters considered in this study, the variance of $r$ is vastly dominated by the first term in that equation. By contrast, the other terms in Eq. (43) are small enough to be ignored. Equation (44) reveals that an increasing slope $F_{h}$ decreases the variance of $\bar{h}$ in an inhibitory network ($w<0$). Intuitively, this effect can be understood by the increase of the effective leakage $(1-wF_{h})/\tau$ in the linearized dynamics of $\bar{h}$ which dampens the input fluctuations. In contrast, the variance of the population rate can increase with increasing slope because the quadratic prefactor $F_{h}^{2}$ can overcompensate the decrease of $\text{var}(\bar{h})$. To make this argument more clear in the theoretical discussion, we further assume that $wF_{h}\gg 1$ (which holds for our parameter setting). Under this assumption, the variance of the population-averaged input potentials becomes inversely proportional to $F_{h}$,
$\text{var}(\bar{h})\approx\left(\frac{\mu_{0}}{2\tau N}+\frac{\sigma_{\text{ext}}^{2}}{-2w}\right)\frac{1}{F_{h}}.$ (45)
In turn, the quadratic factor $F_{h}^{2}$ due to the conversion from variance of $\bar{h}$ to variance of $r$ leads effectively to a linear dependence of $\text{var}(r)$ on $F_{h}$,
@@
The variance formulas (45) and (46) share a common factor that represents the sum of two contributions to the variability. The first term of the sum corresponds to the intrinsically generated variability caused by finite-size fluctuations of order $\sqrt{r_{0}/N}$ in the mesoscopic MF dynamics, Eq. (20). We recall that the dependence of these fluctuations on the firing rate $r_{0}=-\mu_{0}/w$, arises from the Poisson spiking noise at the microscopic level. Thus, the Poisson property that the variance scales proportionally to the mean causes finite-size fluctuations that increase with the mean stimulus strength $\mu_{0}$ (see below). We emphasize that this contribution to the variability is a clear finite-size effect that vanishes in the limit $N\to\infty$. In contrast, the second term of the sum in Eqs. (45) and (46) represents externally generated variability proportional to the variance of the common external stimulus.
Figure 6: Variance of the stochastic population rate $r(t)$ for varying mean external stimulus $\mu_{0}$. (a) Moderate-size network of $N=10^{3}$ neurons whose inputs exhibit internally generated, finite-size noise but no external noise, $\sigma_{\text{ext}}=0$. Different in-degree $C$ are encoded by color. Left: Response to a step stimulus, where $\mu_{0}$ jumps from $28$ mV to 5050 mV at time $t=0.04$ s. Blue/orange line: simulations of MF2 for $C=100$ ($p=0.1$) and $C=1000$ ($p=1$), the latter ($p=1$) equivalent to MF1 or the mean-connectivity approximation of the former ($p=0.1$). Black dotted: corresponding theory for the stationary variances, Eq. (43). Gray dashed: stimulus onset. Inset: magnified data for $C=100$. Symbols represent stationary variances of network simulations and theory, see description of right panel. Right: Stationary variance of $r(t)$ vs. mean stimulus strength $\mu_{0}$. Symbols: simulations for $C=100$ of quenched network (green circles), MF1 (orange triangles), and MF2 (blue diamonds). Solid lines: full MF1 (orange) and MF2 theory (blue). Black dashed: corresponding analytical approximations, Eq. (47); dash-dotted: MFsparselim. The gray dashed line marks the border of the region of interest $\mu_{0}\leq-wr_{\text{m}}/2$ where the hazard function at the fixed point, $\phi(\bar{h}_{0})$ is convex. (b) Analogous to (a) for a much larger network ($N=5\cdot 10^{4}$) with negligible internally generated finite-size noise but with additional external noise $\sigma_{\text{ext}}^{2}=1$ mV2. The connection probability is varied by different values of $C$ (encoded by color). Orange line: common MF1 theory or mean-connectivity network. Dash-dotted lines: MFsparselim. (c) Slope $\hat{F}_{h}$ vs. mean stimulus $\hat{\mu}_{0}$. Dotted: $p=1$ corresponding to the function $Q$, solid line: $p=0.7$. Gray dashed as in (a),(b) right. (d) Same for the product $\hat{\mu}_{0}\hat{F}_{h}$. (c), (d): $N=1000$, $\hat{\beta}=500$ (equivalent to $\beta=5$ mV-1 in (a), (b)).
@@
In the second scenario with constant $C$ and varying population size $N=C/p$, we observe different dependencies of the variances on $p$ compared to the first scenario with constant $N$ (Fig. 5b). In simulations of microscopic networks and 2nd-order MF dynamics, the rate variance increases supralinearly with $p$ and the variance of the mean input potential $\bar{h}$ exhibits a non-monotonic behavior with a maximum inside the range $0<p<1$. Again, the 1st-order MF theory does not capture these behaviors. Note that there is a good qualitative agreement between 2nd-order MF theory and microscopic network simulations. As in the first scenario, however, the variance of the average input potential in the quenched network is underestimated by the annealed network, which, in turn, is quantitatively well matched by the 2nd-order MF theory.
@@
(cf. Eq. (37)). The expression for the rate variance in Eq. (47) is a product of two monotonously increasing functions of $p$ – a linear function (the prefactor of $\hat{F}_{h}$) and a strictly convex function ($\hat{F}_{h}$). Hence, the variance of the population rate is a monotonously increasing, strictly convex function of $p$ explaining our observation in Fig. 5b (left). In contrast, the variance of the mean input potential, Eq. (45), can be rewritten for constant $C$ and dimensionless parameters as